FIND AVG VALUE OF MODULES E POWER X-1 FIND AT (-1,1)

Grade:6

1 Answers

Sher Mohammad IIT Delhi
10 years ago
average value of exp|x-1| between -1 and 1
exp|x-1| = exp(x-1) {for x>1
= exp(1-x){ for x<1


average =(1/2)?exp(x-1) (between -1 and 1)

=(1/2)[exp(x-1)](between -1 and 1)

=(1/2)(exp(1-1)-exp(-1-))=(1/2)(1-exp(-2))